OpenStudy (anonymous):

The equation cos(pi/6-x) is equal to______. A.) 1/2(cosx+ sqrt3 sinx) B.)1/2(cosx- sqrt3 sinx) C.)1/2(sqrt3 cosx- sinx) D.)1/2(sqrt3 cosx+sinx)

OpenStudy (tkhunny):

Have you considered the cosine of a difference formula? \(\cos(\alpha - \beta) = \cos(\alpha)\cos(\beta)+\sin(\beta)\sin(\alpha)\)

OpenStudy (anonymous):

Is the answer C?

OpenStudy (tkhunny):

You tell me. How did you get that?

OpenStudy (anonymous):

I messed up with the signs it is D because it is subtraction first and with the negative in the problem makes it addition

OpenStudy (tkhunny):

Do you believe it?

OpenStudy (anonymous):

yep!

OpenStudy (tkhunny):

That's good enough for me. I won't even check it!